# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

### v1

@@ Appendix E Applications on Real-time LLM Alignment
Following the conformal alignment framework Gui et al. [2024], we randomly split $\mathcal{D}$ into two subsets: a training set $\mathcal{D}_{\mathcal{T}}$ and a calibration set $\mathcal{D}_{\mathcal{C}}$. Using $\mathcal{D}_{\mathcal{T}}$, we train an alignment predictor $g({\bf X})$ to estimate the alignment score based on features of LLM outputs ${\bf X}_{t}$ and compute the predicted alignment scores $\widehat{A}_{t}=g({\bf X}_{t})$ for each $t\in[T]$. Then applying OCTF with $\mathcal{D}_{\mathcal{C}}$ to select new images whose generated reports are aligned with expert standards ensures finite-sample online FDR control according to Theorem 3. The conformal $p$-value for $t\in[T]$ is:
$p_{t}=\frac{1+\sum_{i\in{\mathcal{C}}_{0t}}{\mathbb{I}}\{\widehat{A}_{i}\geq\widehat{A}_{t}\}}{1+|{\mathcal{C}}_{0t}|},$ (33)
@@
4: Initialize ${\mathcal{C}}_{0t}=\{i\in{\mathcal{C}}:Y_{i}\leq c\}$
@@
Appendix F Additional Experiments Results
@@
Specifically, we report additional results for the classification example (Scenarios IV) as well as for a regression example (Scenario VI). The corresponding data generation process for Scenario VI is detailed below:
  • Scenario VI (Regression example): $Y=-0.5X_{1}^{2}+\exp{X_{2}}+(X_{3}+X_{4})^{2}+\varepsilon,$ with ${\bf X}\sim\mathcal{N}_{4}({\bf 0},\mathbf{I}_{4})$ and $\varepsilon\sim\mathcal{N}(0,2)$. The target region is $\mathcal{A}=[c,\infty)$, where $c$ is the $1-\pi_{1}$ quantile of $Y$.
@@
Results.
The results for Scenario VI using a fixed training algorithm (random forest) are shown in Figure 8. The performance trends are similar to those observed in Scenario IV.
Figure 8: Results for Scenario VI: values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportions $\pi_{1}$. The black dashed line denotes the FDR level $\alpha=0.2$.
The results for Scenarios IV and VI under different training algorithms—RF, SVM, and NN—with varying initial calibration sizes are presented in Figure 9-10. Thanks to the online updating of the calibration dataset, even a small initial calibration size does not significantly impact performance. While the choice of predictive model $\widehat{\mu}$ does affect power, all methods maintain valid FDR control. Notably, our SF and LF methods consistently outperform the baselines across all models, benefiting from the distribution-free and model-agnostic nature of online conformal $p$-values. The variation in performance across different algorithms further underscores the importance of careful model selection in practice.
@@ Results.
Figure 10: Results for Scenario VI: $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ vs. initial calibration size $n$ ($\pi_{1}=0.5$, $\alpha=0.2$).
The results with model selection for Scenarios IV and VI are shown below Figure 11-Figure 12. In both settings, the performance gap between the Opt methods and their randomly selected counterparts is also pronounced, with Opt-SF once again achieving the highest power.
@@
Figure 12: Results for Scenario VI: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportion $\pi_{1}$. The black dashed lines denote the FDR level $\alpha=0.1$.

### v2

@@ Appendix D Applications on Real-time LLM Alignment
Following the conformal alignment framework Gui et al. [2024], we randomly split $\mathcal{D}$ into two subsets: a training set $\mathcal{D}_{\mathcal{T}}$ and an initial calibration set $\mathcal{D}_{{\mathcal{C}}}$. Using $\mathcal{D}_{\mathcal{T}}$, we train an alignment predictor $g({\bf X})$ to estimate the alignment score based on features of LLM outputs ${\bf X}_{t}$ and compute the predicted alignment scores $\widehat{A}_{t}=g({\bf X}_{t})$ for each $t\in[T]$. Then applying OCTF with $\mathcal{D}_{{\mathcal{C}}_{0t}}$ to select new images whose generated reports are aligned with expert standards ensures finite-sample online FDR control according to Theorem 4, where the calibration set ${\mathcal{C}}_{0t}$ is updated online as the feedback revealed. The conformal $p$-value for $t\in[T]$ is:
$p_{t}=\frac{1+\sum_{i\in{\mathcal{C}}_{0t}}{\mathbb{I}}\{\widehat{A}_{i}\geq\widehat{A}_{t}\}}{1+|{\mathcal{C}}_{0t}|}.$ (29)
@@
4: Initialize ${\mathcal{C}}_{0t}=\{i\in{\mathcal{C}}:A_{i}\leq c\}$
@@
Appendix E Additional Experiments Results
@@
Specifically, Appendix E.1 reports additional results for a regression example (Scenario V); Appendix E.2 reports results under different training algorithms for $\hat{\mu}$; Appendix E.3 reports additional results after score selection; Appendix E.4 reports empirical mFDR results; and Appendix E.5 reports results on auxiliary calibration for score selection.
E.1 Results on synthetic data for a regression task
The corresponding data generation process for the regression task is detailed below:
  • Scenario V (Regression example): $Y=-0.5X_{1}^{2}+\exp{X_{2}}+(X_{3}+X_{4})^{2}+\varepsilon,$ with ${\bf X}\sim\mathcal{N}_{4}({\bf 0},\mathbf{I}_{4})$ and $\varepsilon\sim\mathcal{N}(0,2)$. The target region is $\mathcal{A}=[c,\infty)$, where $c$ is the $1-\pi_{1}$ quantile of $Y$.
@@ E.1 Results on synthetic data for a regression task
The results for the regression example (Scenario V) using a fixed training algorithm (random forest) are shown in Figure 14. The performance trends are similar to those observed in the classification case (Scenario IV).
Figure 14: Results for Scenario V: values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportions $\pi_{1}$. The black dashed line denotes the FDR level $\alpha=0.2$.
E.2 Additional experiments for different training algorithms
The results for Scenarios IV and V under different training algorithms—RF, SVM, and NN—with varying initial calibration sizes are presented in Figure 15-16. Thanks to the online updating of the calibration dataset, even a small initial calibration size does not significantly impact performance. While the choice of predictive model $\widehat{\mu}$ does affect power, all methods maintain valid FDR control. Notably, our SF and LF methods consistently outperform the baselines across all score functions, benefiting from the distribution-free and model-agnostic nature of online conformal $p$-values. The variation in performance across different algorithms further underscores the importance of careful score selection in practice.
@@ E.2 Additional experiments for different training algorithms
Figure 16: Results for Scenario V: $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ vs. initial calibration size $n$ ($\pi_{1}=0.5$, $\alpha=0.2$).
E.3 Additional experiments on score selection
The results with score selection for Scenarios IV and V are shown below Figure 17-Figure 18. In both settings, the performance gap between the Opt methods and their randomly selected counterparts is also pronounced, with Opt-SF once again achieving the highest power.
@@ E.3 Additional experiments on score selection
Figure 18: Results for Scenario V: the values of $\operatorname{FDR}(T)$ and $\operatorname{Power}(T)$ at stopping time $T$ across different non-null proportion $\pi_{1}$. The black dashed lines denote the FDR level $\alpha=0.1$.
E.4 Additional experiments results on empirical mFDR